# Unified Medical Image Pre-training in Language-Guided Common Semantic Space

### v1

@@ Abstract
Vision-Language Pre-training (VLP) has shown the merits of analysing medical images, by leveraging the semantic congruence between medical images and their corresponding reports. It efficiently learns visual representations, which in turn facilitates enhanced analysis and interpretation of intricate imaging data. However, such observation is predominantly justified on single-modality data (mostly 2D images like X-rays), adapting VLP to learning unified representations for medical images in real scenario remains an open challenge. This arises from medical images often encompass a variety of modalities, especially modalities with different various number of dimensions (e.g., 3D images like Computed Tomography). To overcome the aforementioned challenges, we propose an Unified Medical Image Pre-training framework, namely UniMedI, which utilizes diagnostic reports as common semantic space to create unified representations for diverse modalities of medical images (especially for 2D and 3D images). Under the text’s guidance, we effectively uncover visual modality information, identifying the affected areas in 2D X-rays and slices containing lesion in sophisticated 3D CT scans, ultimately enhancing the consistency across various medical imaging modalities. To demonstrate the effectiveness and versatility of UniMedI, we evaluate its performance on both 2D and 3D images across 10 different datasets, covering a wide range of medical image tasks such as classification, segmentation, and retrieval. UniMedI has demonstrated superior performance in downstream tasks, showcasing its effectiveness in establishing a universal medical visual representation.
@@ 1 Introduction
Figure 1: An example showing X-ray (up) and CT scan (down) both demonstrate similar abnormality, recording in the report.
@@
In recent years, the field of medical image analysis has witnessed significant advancements, largely driven by the application of deep learning techniques and the increasing availability of medical imaging data. Notably, Visual-Language Pre-training (VLP) (Huang et al., 2021; Boecking et al., 2022; Bannur et al., 2023) attracts lots of attention, as it reduces the need for costly and time-consuming manual annotations by leveraging the vast amount of information in radiology reports and unlabeled data. Despite these success, further expanding the data scale for medical VLP remains non-trivial, because the availability of single-modality medical images is limited, especially when compared to the general domain. This introduces a strong need to integrate multi-modality medical images (e.g., X-rays, Computed Tomography (CT) and Magnetic Resonance Imaging(MRI)) within a unified VL framework. However, fully leveraging the information across multi-modal images within this VL framework is unexplored.
@@
On the above aspect, the inherent heterogeneity of medical imaging from different modalities obstructs their effective integration. One obvious and important problem is that medical images have different dimensions. For example, X-rays are 2D images, while CT scans are 3D images. To tackle this challenge, we start from the following key observation: despite big differences, medical images from various modalities share a common semantic latent space, which captures the underlying features of an individual’s health status, and such status are reflected in medical reports via language. As shown in Fig. 1, the X-ray and CT scan can contribute to a comprehensive understanding of pneumonia, reflecting the commonality within the latent space, and these abnormalities are listed in reports. This observation motivate us to map data from various medical image modalities into the shared semantic space, which is guided by language in reports. This strategy not only tackles data-related issues but also fosters synergy and collaboration among distinct modalities, ultimately resulting in a more holistic understanding of an individual’s health condition.
@@
However, creating a unified model that effectively maps data from different sources into a common space for combined learning is challenging, even with language guidance in reports. Figure 1(a) demonstrates the representation space of two distinct modalities with different dimensions (i.e., 2D X-rays and 3D CT scans) when trained individually via VLP. They are far apart in the representation space, even with same pathological information in reports. Furthermore, Figure 1(b) shows simply unifying them in one model does not solve the problem. Although the distance between representations of two modalities are shortened to some extent, their representations remain insufficiently compact, since only little space are shared between them.
@@
Figure 2: t-SNE visualizations of image representations by models trained with different methods (2D: X-rays, 3D: CT, both modalities denote the same disease, pneumonia.). (a) Two models for different image modalities are trained individually in separate VLP process. (b) One models for different image modalities are trained in one VLP processes, but without designes in UniMedI. (c) UniMedI. Learning a common semantic space for different medical images is non-trivial, even with language guidance, and UniMedI can well handle this integration. We use circles to highlight differences between different images.
@@
To address the above challenge, we propose UniMedI, a novel Unified VL framework, designed to effectively integrate Medical multi-modal Images into a language-guided common semantic space. First, under the dilemma that paired 2D and 3D medical images are unavailable, and naively integration is not effectively as we shown above, we first design an attentive selection method to accurately identify text-relevant 2D slices without extra annotations. This builds a data bridge between 2D and 3D medical images. Then, we devise a cross-dimensional VLP method to bring both 3D data and selected 2D slices closer to the same report representation space, constructing a unified VL framework. Moreover, we introduce a self-distillation technique using a teacher-student structure and construct a masking and recovery task, further enhancing the associations between 2D and 3D data within the image space. Figure 1(c) shows UniMedI significantly reduces the distance between 2D and 3D features after undergoing our effective design for cross-dimensional pre-training.
@@
To further demonstrate the effectiveness of our approach, we conduct extensive visualizations and experiments to showcase the working mechanisms and superior representational capabilities of our model. We evaluate our UniMedI framework on 10 real-world medical datasets and various downstream tasks (i.e., classification, segmentation and retrieval). The results consistently show superior performance, regardless of whether UniMedI is applied to full-scale data or limited data scenarios. We also provide visualizations on regions and slices selected by UniMedI, verifying our claim that UniMedI can identify key information from both 2D and 3D medical images.
@@ 2 Related Work
Medical Self-supervised Learning
In the domain of medical image analysis, a number of self-supervised learning (SSL) techniques have been developed to exploit the unique characteristics of medical data. These methods construct feature embedding spaces by designing pre-text tasks, such as solving jigsaw puzzles Noroozi & Favaro and inpainting tasks Pathak et al. (2016). Recently, researchers have explored the use of 3D convolutional neural network (CNN) architectures while retaining established SSL tasks on 2D CNNs Tang et al. (2022). However, the diversity of medical data poses a significant challenge, as the development of a unified visual representation that adequately captures the intricacies of different data types remains a crucial yet complex task that requires further investigation. To address this challenge, Xie et al. (2022) proposed Unimiss, a universal medical self-supervised representation learning framework that overcomes the dimensionality barrier. Furthermore, Nguyen et al. (2023) introduced Joint, an SSL framework capable of accommodating various data dimensions and generating versatile pre-trained weights for both 2D and 3D downstream applications. These approaches have made notable contributions to handling data from different modalities. However, they have given relatively less attention to the relationships and connections between different types of medical data.
Medical Vision-Language Processing
Medical Vision-Language Processing (VLP) has emerged as a promising approach for learning medical visual representations by leveraging naturally occurring paired descriptive text Zhang et al. (2022). Huang et al. (2021) propose Gloria, an attention-based framework that contrasts image sub-regions and words in the paired report to learn global and local representations. Wang et al. (2022) further optimize the framework from the perspective of disease in their method MGCA. These methods exhibit remarkable performance in various downstream tasks involving medical images. However, the application of medical VLP is primarily limited to 2D images, mainly due to the limited availability of extensive 3D medical image-text datasets. Compared to 2D medical image-text pairs, 3D images and reports contain more abundant information, which offers clear advantages for learning visual representations. While some methods Liu et al. (2023); Chen et al. (2023) attempt to address this limitation by converting 3D data into 2D slices and subsequently employing generative models to generate captions for 3D medical data, this approach results in a loss of the original 3D volume structure information. Therefore, it is imperative to develop strategies that can effectively harness the valuable information present in 3D images and reports while preserving the structural integrity of the data. This will facilitate the enhancement of the learning process for visual representations in medical VLP.
@@ 3 Methodology
Figure 3 illustrates UniMedI and its designs to realize integration of 2D and 3D medical images. Generally, to overcome the challenges that no paired 2D and 3D image data exists, UniMedI employes the following pipeline. When the input is a 3D volume, we first extract a portion of 2D slices from it which most relevant to the report, and then regard the selected slices as 2D image. Those selected 2D slices are fed into the network along with the original 3D volume, allowing us to jointly learn the relationships between 2D, 3D, and radiology reports, and ultimately form a unified feature space. When the input is a 2D image, the slice selection process is omitted.
In Section 3.1, we demonstrate our designed attentive slice selection method, which can identify more relevant 2D slices in 3D data related to the report text, helping us learn the unified space between 2D and 3D data guided by report. In Section 3.2, we design a method to bring together 3D data and selected 2D slices closer to the same report representation, which serves as the foundation for our language-guided construction of a unified model. In Section 3.3, we design a self-distillation technique to EMA teacher for the visual encoder, constructing image-level and patch-level contrastive learning tasks, further enhancing the connection between 2D and 3D data.
Figure 3: Illustration of the proposed UniMedI framework. To effectively integrate multi-modal medical images in a language-guided common semantic space, UniMedI incorporates following designs. (1) An strategy of attentive slice selecting from 3D volume to bridge 2D and 3D images even without paired 2D and 3D data (details shown in Fig. 4). The concatenated inputs of 2D and 3D allow us to perform joint modeling across dimensions. (2) Shared backbone $E_{v}$ for 2D and 3D images and separate tokenizer $T_{2D}$ and $T_{3D}$. (3) Language-guidance for unified image representation provided by the language encoder $E_{l}$ and vision-language loss $L_{vl}$. (4) Self-distillation (implemented by image contrastive learning loss $L_{icl}$ and patch contrastive learning loss $L_{pcl}$) to enhance interactions between images tokens from different modalities. The distillation target comes the teacher network $\overline{E}_{v}$, which is updated by exponential moving averaged (EMA) over the student network $E_{v}$.
3.1 Attentive slice selection
@@
In order to construct a cross-modal unified representation space, we have chosen language as the bridge. Therefore, we need to extract key information from various image modalities that correspond to the information in medical reports. Particularly, important 2D slices relevant with reports should be selected from 3D volume. This process is similar to how doctors view CT scans; they also base their report descriptions on some important slices.
As shown in Figure 4, in order to better locate the lesion-related 2D slices in the 3D data, we use the attention weights of the $\left[CLS\right]$ token in the EMA teacher as the basis for calculation. The visual encoder’s $\left[CLS\right]$ token is directly supervised by the radiology report features from the language encoder, reflecting the most likely lesion areas described in the report. For the attentive score at token location $P$:
@@
$s^{P}=\frac{1}{HL}\sum_{l=1}^{L}\sum_{h=1}^{H}\operatorname{Softmax}\left(\frac{\mathbf{f}^{q}_{lh}({CLS})\cdot{{\mathbf{f}^{k}_{lh}(P)}}}{\sqrt{C}}\right),$ (1)

### v2

@@ Abstract
Vision-Language Pre-training (VLP) has shown the merits of analysing medical images. It efficiently learns visual representations by leveraging supervisions in their corresponding reports, and in turn facilitates analysis and interpretation of intricate imaging data. However, such observation is predominantly justified on single-modality data (mostly 2D images like X-rays), adapting VLP to learning unified representations for medical images in real scenario remains an open challenge. This arises from medical images often encompass a variety of modalities, especially modalities with different dimensions (e.g., 3D images like Computed Tomography), and there are almost no paired multi-dimension data here. To overcome the aforementioned challenges, we propose an Unified Medical Image Pre-training framework, namely UniMedI, which utilizes diagnostic reports as common semantic space to create unified representations for diverse modalities of medical images (especially for 2D and 3D images). Under the text’s guidance, UniMedI effectively select text-related 2D slices from sophisticated 3D volume, which acts as pseudo-pairs to bridge 2D and 3D data, ultimately enhancing the consistency across various medical imaging modalities. To demonstrate the effectiveness and versatility of UniMedI, we evaluate its performance on both 2D and 3D images across several different datasets, covering a wide range of medical image tasks such as classification, segmentation, and retrieval. UniMedI has demonstrated superior performance in downstream tasks, showcasing its effectiveness in establishing a universal medical visual representation.
Keywords:
Vision-Language Pre-Training Medical Image Multi-Modality
@@ 1 Introduction
In recent years, deep models with Visual-Language Pre-training (VLP) [14, 4, 3] attract lots of attention in medical image analysis, as VLP reduces the need for costly and time-consuming manual annotations by leveraging the vast amount of information in radiology reports and unlabeled data. Despite these success, further expanding the data scale for medical VLP remains non-trivial, because the availability of single-modality medical images is limited, especially when compared to the general domain. This introduces a strong need to integrate multi-modality medical images (e.g., X-rays, Computed Tomography (CT) and Magnetic Resonance Imaging (MRI)) within a unified Vision-Language (VL) framework. However, fully leveraging the information across multi-modal images within this VL framework is unexplored.
(a) An example showing X-ray (up) and CT scan (down) both demonstrate similar abnormality, recording in the reports (right).
(b) t-SNE visualization of report representations by a pre-trained BioClinicalBERT [1]. Implementation details are in Appendix A.
Figure 1: Observations that motivate the language-guided strategy for integrating 2D and 3D medical images in VLP.
@@
On the above aspect, the inherent heterogeneity of medical imaging from different modalities obstructs their effective collaboration and integration. One obvious and important problem is that medical images have different dimensions. For example, X-rays are 2D images, while CT scans are 3D images. Nevertheless, when looking at their reports, which provide supervisions for the visual representation learning in VLP, it can be noticed that medical reports of images with different dimensions generally reflects same contents, which are an individual’s health status. As shown in Fig. 1(a), two reports mainly list the same abnormalities, and the X-ray and CT scan can contribute to a comprehensive understanding of pneumonia. The rationale behind is that despite big differences, medical images from various modalities are all used for capturing the underlying features of an individual’s health status, and such status are reflected in medical reports via language. Furthermore, we visualize the representation space of reports of 2D and 3D images in Fig. 1(b), showing that their representations lie together in a common semantic space. This observation motivate us to map data from various medical image modalities into the shared semantic space, which is guided by language in reports. This strategy not only tackles data-related issues but also fosters synergy and collaboration among distinct modalities, ultimately resulting in a more holistic understanding of an individual’s health condition.
@@
However, under the dilemma that paired 2D and 3D medical images with reports are unavailable, creating a unified model that effectively maps data of different modalities into a common space for combined learning is challenging, even with language guidance in reports. Fig. 2(a) demonstrates the representation space of two distinct modalities with different dimensions (i.e., 2D X-rays and 3D CT scans) when trained individually via VLPs. They are far apart in the representation space, even with same pathological information in reports. Furthermore, Fig. 2(b) shows simply unifying them in one model does not solve the problem. Although the distance between representations of two modalities are shortened to some extent, the degree of fusion is still far lower than that of report representations, which provides supervisions in VLP.
@@
Figure 2: t-SNE visualizations of image representations by models trained with different VLP methods (2D: X-rays, 3D: CT). Both modalities are on the same disease, i.e., pneumonia, and the differences among sub-figures are highlighted with circles. (a) Two models for different image modalities are trained individually in separate VLP process. (b) One models for different image modalities are trained in one VLP processes, but without designs in UniMedI. (c) UniMedI, introducing pseudo-paired ‘2D’ and 3D images in one unified framework.
@@
To address the above challenge, we propose UniMedI, a novel Unified VL framework, designed to effectively integrate Medical multi-modal Images into a language-guided common semantic space, making them better help with each other. Notably, UniMedI introduces a technique for creating the pseudo-pairs, enabling inputting both ‘2D’ and 3D images in one forward pass. The pseudo-pair is created via an attentive selection method that can accurately identify text-relevant 2D slices without extra annotations. This builds a data bridge between 2D and 3D medical images. Besides, our unified VL framework contains one model and one VLP for different dimensional image data, bringing both 3D data and selected 2D slices closer to the same report representation space. Moreover, a self-distillation technique is employed for the auxiliary masking and recovering task, further enhancing the interactions between 2D and 3D data within the image space. Fig. 2(c) shows UniMedI significantly reduces the distance between 2D and 3D features after undergoing our effective designs for cross-dimension pre-training, making it more like the report representation space. Our experiment results (see Fig. 6(d)) also verify that UniMedI can make 2D and 3D images help with each other, making most of them in the pre-training stage.
@@
We conduct extensive experiments on several real-world medical datasets and various downstream tasks (i.e., classification, segmentation and retrieval). The consistently superior performance well demonstrates the effectiveness and capacity of our UniMedI framework, regardless of whether it is applied to full-scale data or limited data scenarios. We also provide visualizations on slices selected by UniMedI, verifying our claim that UniMedI can identify key information from 3D medical images and create pseudo-pairs for integration.
@@ 2 Related Work
2.0.1 Medical Self-Supervised Learning
In the domain of medical image analysis, a number of self-Supervised learning (SSL) techniques have been developed to exploit the unique characteristics of medical data. These methods construct feature embedding spaces by designing pre-text tasks, such as solving jigsaw puzzles [26] and inpainting tasks [27]. Recently, researchers have explored the use of 3D convolutional neural network (CNN) architectures while retaining established SSL tasks on 2D CNNs [33]. However, the diversity of medical data poses a significant challenge, as the development of a unified visual representation that adequately captures the intricacies of different data types remains a crucial yet complex task that requires further investigation. To address this challenge, [36] proposed UniMiSS, a universal medical self-supervised representation learning framework that overcomes the dimensionality barrier. Furthermore, [25] introduced Joint, an SSL framework capable of accommodating various data dimensions and generating versatile pre-trained weights for both 2D and 3D downstream applications. These approaches have made notable contributions to handling data from different modalities. However, they have given relatively less attention to the relationships and connections between different types of medical data.
2.0.2 Medical Vision-Language Pre-Training
Recently, Medical Vision-Language Pre-Training (VLP) has emerged as a promising approach for learning medical visual representations by leveraging naturally occurring paired descriptive text [40]. [14] propose Gloria, an attention-based framework that contrasts image sub-regions and words in the paired report to learn global and local representations. [34] further optimize the framework from the perspective of disease in their method MGCA. These methods exhibit remarkable performance in various downstream tasks involving medical images. However, the application of medical VLP is primarily limited to 2D images, mainly due to the limited availability of extensive 3D medical image-text datasets. Compared to 2D medical image-text pairs, 3D images and reports contain more abundant information, which offers clear advantages for learning visual representations. While some methods [21, 6] attempt to address this limitation by converting 3D data into 2D slices and subsequently employing generative models to generate captions for 3D medical data, this approach results in a loss of the original 3D volume structure information. Therefore, it is imperative to develop strategies that can effectively harness the valuable information present in 3D images and reports while preserving the structural integrity of the data. This will facilitate the enhancement of the learning process for visual representations in medical VLP.
@@ 3 Methodology
To effectively integrate multi-modal medical images into a language-guided common semantic space, we propose UniMedI framework and present its components in this section. We first show the overall pipeline and outline the key designs in Section 3.1, and then explain details of each design in the following sub-sections.
3.1 Pipeline and Key Designs
The illustration of UniMedI is in Fig. 3. Generally, UniMedI is a vision-language pre-training framework, in which medical images and their reports are encoded by two encoders, separately, and they are jointly learned via the VL contrastive learning. The distinction of UniMedI is that it can effectively taking both 2D and 3D images in an unified way, tackling data scarcity issue in the medical domain and making them help with each other.
First, to overcome the challenges that no paired 2D and 3D image data exist, UniMedI introduces a novel attentive slice selection strategy for creating pseudo-pairs. Here, when the input is a 3D volume, a portion of 2D slices which most relevant to the report are extracted from it, and then the selected slices are regarded as 2D image, thus forming pseudo 2D-3D pairs. Those selected 2D slices are fed into the network together with the original 3D volume in a forward pass, allowing jointly learning the relationships between 2D, 3D, and radiology reports, and ultimately form a unified feature space. When the input is a 2D image, the slice selection process is omitted. After that, one vision encoder maps all multi-modal images, including original 2D and 3D ones and also selected 2D slices into the representation space. This vision encoder has tokenizers $T_{2D}$ and $T_{3D}$ for 2D and 3D images, respectively, and a shared backbone $E_{v}$ for enabling better integration. Then, the model which contains the vision encoder and text encoder $E_{l}$ are learned end-to-end in one VLP via contrastive learning loss $L_{vl}$. Thus, 2D and 3D images can all be encoded into a common semantic space supervised by language information in reports. Finally, UniMedI further introduces an auxiliary masking and recovering task for enhancing interactions among 2D and 3D tokens. This is completed via the self-distillation technique.
Our proposed attentive slice selection method is demonstrated in Section 3.2. The VL contrastive learning used in our framework is explained in Section 3.3. The self-distillation technique is elaborated in Section 3.4.
Figure 3: Illustration for the proposed UniMedI framework. The overall pipeline is shown in the left part, and key designs are displayed in the right part.
3.2 Creating Pseudo-Pairs via Attentive Slice Selection
@@
According to our observation, we choose language in reports as the bridge to construct a cross-modal unified representation space. Therefore, our goal is to select 2D slices that relevant most with reports from 3D volume. Note that this process is similar to how doctors view CT scans. They also base their report descriptions on some important slices.
To achieve the goal, we design a novel attentive slice selection strategy which is guided by language, and show it in Fig. 4. Basically, the relevance scores to the report are calculated for each slice and top $k$ slices with highest scores are selected. Here, the attention value between the $\left[CLS\right]$ token produced by the vision encoder and other visual tokens are used as the basis for calculating relevance scores. As common VL pre-training methods [28, 9, 38, 19], this $\left[CLS\right]$ token is directly supervised by language, reflecting the most important lesion information in reports. Therefore, this token is likely to have large attention values with lesion-related 2D slices, and thus can be used for selection. Formally, the attention value vector $\mathbf{v}\in\mathbb{R}^{P}$ between $\left[CLS\right]$ token and $P$ image patch tokens is denoted as:
@@
$\mathbf{v}=\frac{1}{HL}\sum_{l=1}^{L}\sum_{h=1}^{H}\operatorname{Softmax}\left(\frac{\mathbf{q}_{lh}({\left[CLS\right]}){{\mathbf{K}_{lh}(\left[Patch\right])}}}{\sqrt{C}}\right).$ (1)